# RAMAC: Multimodal Risk-Aware Offline Reinforcement Learning and the Role of Behavior Regularization

### v1

@@ D.2 Extended Stochastic-D4RL Results
At the fixed 1000–step evaluation, the ranking patterns largely match the main text, but the mechanisms are task–dependent. Flow–based policies (FlowQL/RAFMAC) often reach higher mean by 1000 steps because flow matching uses a deterministic ODE with a short generative path and low-variance policy gradients; combined with velocity-matching BC, this yields fast on-manifold improvement. CVaR outcomes depend on lower-tail calibration of the distributional critic: with smooth, non-terminating penalties (e.g., HalfCheetah) RADAC/RAFMAC already improve CVaR at 1000 steps, whereas with sparse, terminating hazards (e.g., Hopper) ORAAC’s anchor regularization provides more stable early CVaR and mean. Walker2d sits in between: RAFMAC attains the highest mean at 1000 steps, and CVaR leadership alternates between FlowQL and RAFMAC depending on the dataset variant.
@@
Across datasets, RADAC populates the upper-left region of the frontier: for comparable return, it tends to incur fewer violations. Only for halfcheetah–medium–expert, RADAC sometimes drifts up-right (higher return with slightly more violations) because the penalty is light and non-terminating, so near-threshold speed pays off, consistent with its best Mean/CVaR. Mechanistically, diffusion with CVaR guidance enables fine-grained reweighting away from safety thresholds while BC keeps samples on-manifold, so trajectories in the plot drift left (fewer violations) without sacrificing return. RAFMAC pushes the top of the frontier in mean—most clearly on walker2d and halfcheetah-m-r—and is competitive in CVaR (Table 4). Deterministic ODE transport with low-variance policy gradients and velocity-matching BC yields fast on-manifold improvement; in the Pareto view this appears as high-return points with modest violation counts. Because the transport is geometry-preserving, boundary-adjacent mass tends to thin rather than disappear abruptly; CVaR improves as the velocity field adapts. ORAAC forms the frontier on hopper-m-e with few violations and strong returns, matching its leading scores under terminating pose hazards. In other settings it remains reliably conservative (low violations) at the cost of mean on some tasks, consistent with anchor-based regularization. FlowQL often achieves high-mean points but with comparatively higher violation counts in the Pareto plot. Without tail-aware guidance, safety depends on the expected-value critic and task smoothness, explaining the variability across datasets. DiffusionQL exhibits wider scatter: runs either reach moderate returns with elevated violations or collapse to low-return, near-zero violation regions. This variability is consistent with value-only guidance under stochastic penalties and matches its weaker CVaR. CODAC clusters in the low-return/low-violation corner across tasks, as expected from conservative critics.
@@
D.3 Ablation Study
@@
Appendix E Experimental Details
E.1 2D Synthetic Task Details
@@ Appendix E Experimental Details
E.2 Stochastic-D4RL MuJoCo Suite
Datasets We adopt the stochastic MuJoCo protocol for risk-sensitive offline RL, following (Urpí et al., 2021). Policies are evaluated on
$\{\textsc{hopper},\ \textsc{walker2d},\ \textsc{halfcheetah}\}\ \times\ \{\textsc{medium-expert},\ \textsc{medium-replay}\},$
Compared to prior work, we prefer medium-expert and medium-replay to validate both risk sensitivity and policy expressiveness under multimodal action distributions. For training, we relabel per-transition rewards in the offline datasets to inject stochastic hazards (velocity or torso-pitch thresholds with Bernoulli penalties and early termination); the same hazard model is used at evaluation. This ensures the critic and the policy are trained on the risk-aware rewards rather than only being tested under hazards.
@@ E.2 Stochastic-D4RL MuJoCo Suite
E.3 Baselines: Implementation & Hyperparameters
@@
  • CODAC (Ma et al., 2021) (distributional conservative learning). We primarily use the CVaR-optimizing specification (“CODAC-C”, $\mathrm{CVaR}_{0.1}$ objective).
  • ORAAC (Urpí et al., 2021) (offline risk-averse actor–critic). A distributional critic with imitation-regularized policy optimizing a coherent risk objective.
  • CQL (Kumar et al., 2020) (value pessimism). Non-distributional conservative Q-learning baseline.
  • DiffusionQL (Wang et al., 2023) (expressive risk-neutral diffusion policy).
  • FlowQL (Park et al., 2025) (expressive risk-neutral flow-matching policy).
@@
For each of baselines, we run all baselines ourselves and tune the following parameters or adopt authors’ recommended settings, mirroring the practice in Ma et al. (2021); Urpí et al. (2021); Wang et al. (2023); Park et al. (2025); Kumar et al. (2020).
@@
  • FlowQL (Park et al., 2025): we sweep the policy weight $\alpha\in\{1,10,30,100,1000\}$ per task and report the best-performing setting (selection by $\mathrm{CVaR}_{0.1}$ unless noted).
@@
  • DiffusionQL (Wang et al., 2023): we consider $\eta\in\{0.1,0.5,1.0\}$ for BC coefficient . we use authors’ recommended configuration for other parameters without retuning. We also used the best checkpoint of their model on each benchmark by following their protocol.
@@
  • ORAAC (Urpí et al., 2021): use the paper’s recommended configuration (distributional critic, risk level $\alpha=0.1$, anchor/prior regularization) without additional sweeps.
@@
  • CODAC (Ma et al., 2021): use the paper’s tuned settings for D4RL (risk level $\alpha=0.1$) without further tuning.
@@
  • CQL (Kumar et al., 2020): use the standard conservative coefficient and implementation defaults for MuJoCo locomotion.
E.4 Estimating the OOD Visitation Rate with a 1 –NN Detector
@@
At evaluation time, we estimate the fraction of actions produced by a policy that fall outside the empirical action support of the offline dataset. This fraction is reported as the OOD action rate $\varepsilon_{\text{act}}$ in Sec. 5.3. Let $\mathcal{A}_{\mathcal{D}}=\{a_{i}\}_{i=1}^{N}$ be the set of actions in the offline dataset (per task), and let $\|\cdot\|_{2}$ denote the Euclidean norm in action space. In our MuJoCo tasks, actions are already scaled to [1,1][-1,1] per dimension; we therefore use $\ell_{2}$ directly.
For each dataset action $a_{i}$, compute its nearest neighbour among the other dataset actions and the associated distance
$d_{i}\;=\;\min_{j\neq i}\,\|a_{i}-a_{j}\|_{2}.$
Define the robust scale
$\operatorname{medNN}\;=\;\mathrm{median}\{d_{i}\}_{i=1}^{N}.$
We set the OOD threshold to $\tau\,=\,\kappa\cdot\operatorname{medNN}$ with $\kappa{=}3$ unless otherwise stated.
Label evaluation actions
Let $\mathcal{A}_{\text{eval}}=\{a^{(\text{eval})}_{t}\}_{t=1}^{T}$ be all actions emitted across evaluation rollouts (we use 5 seeds $\times$ 10 episodes/seed by default). For each $a^{(\text{eval})}_{t}$, compute its distance to the dataset action set
$d^{(\text{eval})}_{t}\;=\;\min_{i\in\{1,\dots,N\}}\|a^{(\text{eval})}_{t}-a_{i}\|_{2},$
and assign the indicator
$\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})\;=\;\mathbb{I}\!\left\{d^{(\text{eval})}_{t}\;>\;\tau\right\}.$
In practice we compute $d^{(\text{eval})}_{t}$ via a KD–tree built on $\mathcal{A}_{\mathcal{D}}$ (query time $O(\log N)$ in low to moderate dimensions).
We define the OOD action rate as the per–action frequency
$\varepsilon_{\text{act}}\;=\;\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})$.
@@
We report the mean and standard error over 5 seeds. Because episodes may terminate early under the stochastic wrappers, $T$ is the actual number of executed timesteps (not a fixed horizon), which makes the rate comparable across seeds.
Confidence intervals
@@
For a seed–level rate $\hat{\varepsilon}$ with $T$ trials, we use a binomial approximation for the standard error $\mathrm{SE}=\sqrt{\hat{\varepsilon}(1-\hat{\varepsilon})/T}\,$ and report the across–seed mean $\pm$ s.e.

### v2

@@ E.2 Extended Stochastic-D4RL Results
At the fixed 1000-step evaluation, the ranking patterns largely match the main text, but the mechanisms are task–dependent. Flow–based policies (FlowQL/RAFMAC) often reach higher mean by 1000 steps because flow matching uses a deterministic ODE with a short generative path and low-variance policy gradients; combined with velocity-matching BC, this yields fast on-manifold improvement. CVaR outcomes depend on lower-tail calibration of the distributional critic: with smooth, non-terminating penalties (e.g., HalfCheetah) RADAC/RAFMAC already improve CVaR at 1000 steps, whereas with sparse, terminating hazards (e.g., Hopper) ORAAC’s anchor regularization provides more stable early CVaR and mean. Walker2d sits in between: RAFMAC attains the highest mean at 1000 steps, and CVaR leadership alternates between FlowQL and RAFMAC depending on the dataset variant.
@@
Across datasets, RADAC populates the upper-left region of the frontier: for comparable return, it tends to incur fewer violations. Only for HalfCheetah–medium–expert, RADAC sometimes drifts up-right (higher return with slightly more violations) because the penalty is light and non-terminating, so near-threshold speed pays off, consistent with its best Mean/CVaR. Mechanistically, diffusion with CVaR guidance enables fine-grained reweighting away from safety thresholds while BC keeps samples on-manifold, so trajectories in the plot drift left (fewer violations) without sacrificing return. ORAAC forms the frontier on Hopper-m-e with few violations and strong returns, matching its leading scores under terminating pose hazards. In other settings it remains reliably conservative (low violations) at the cost of mean on some tasks, consistent with anchor-based regularization. FlowQL often achieves high-mean points but with comparatively higher violation counts in the Pareto plot. Without tail-aware guidance, safety depends on the expected-value critic and task smoothness, explaining the variability across datasets. DiffusionQL exhibits wider scatter: runs either reach moderate returns with elevated violations or collapse to low-return, near-zero violation regions. This variability is consistent with value-only guidance under stochastic penalties and matches its weaker CVaR. CODAC clusters in the low-return/low-violation corner across tasks, as expected from conservative critics.
@@
E.3 Ablation Study
E.3.1 Choice of risk distortion
@@
E.3.2 Effect of the tail sample size $K$ and $N$ on the CVaR estimator
To clarify the role of the tail sample size $K$ in our IQN-based CVaR estimator, we run an experiment on a subset of D4RL, HalfCheetah-Medium-Replay-v2 and Walker2d-Medium-Replay-v2. During training we randomly sample 2,0002{,}000 states from the replay buffer and store them as a fixed evaluation set $\{s_{j}\}_{j=1}^{2000}$. For each state $s_{j}$ , each $K\in\{2,4,8,16\}$, and each risk level $\alpha\in\{0.05,0.1,0.2\}$ we estimate $\mathrm{CVaR}_{\alpha}$ using the offline-selected RADAC policy (trained with $\alpha_{\text{train}}=0.1$).
Given a state $s$, we draw $K$ i.i.d. pairs $(a_{k},\tau_{k})$ with $a_{k}\sim\pi_{\theta}(\cdot\mid s)$ and $\tau_{k}\sim\mathrm{Unif}(0,\alpha)$, evaluate both distributional critics $Z_{\phi_{1}},Z_{\phi_{2}}$, and form the tail-sampling CVaR estimator
$\widehat{\mathrm{CVaR}}_{\alpha}(s)=\frac{1}{K}\sum_{k=1}^{K}\min_{i\in\{1,2\}}Z_{\phi_{i}}\bigl(s,a_{k};\tau_{k}\bigr).$ (20)
For each $(K,\alpha)$ we repeat this procedure $100$ times per state (using new $(a_{k},\tau_{k})$ draws each time), compute the variance of $\widehat{\mathrm{CVaR}}_{\alpha}(s_{j})$ across the $100$ repetitions for each state $s_{j}$, and then aggregate these per-state variances by their mean and standard deviation over $j=1,\dots,2000$.
Figure 11 reports the resulting estimator variance as a function of $K$ for $\alpha\in\{0.05,0.1,0.2\}$. Across both tasks and all three risk levels the variance decreases approximately at a $1/K$ rate: using only $K=2$ tail samples yields noisy CVaR estimates, while increasing $K$ to $4$ and $8$ substantially reduces the variance. The marginal improvement from $K=8$ to $K=16$ is much smaller, despite doubling the number of tail samples. This supports our choice of a moderate tail size (with $K$ in the range 881616 across tasks in our experiments), which provides a good trade-off between estimator noise and computational cost. Note that the absolute variance scale differs between environments.
Role of the number of quantiles $N$.
The tail estimator in Eq. 20 depends on the number of tail samples $K$, while the IQN critic itself is trained with $N$ quantile samples per state–action pair. In all experiments we use moderate values $N\in\{16,32\}$ (see Table. 6), which are standard in prior IQN-based work and make the critic updates stable. Increasing $N$ primarily reduces the variance of the critic update and smooths the learned value distribution, but once $N$ is in this moderate range we do not observe qualitative changes in the CVaR estimates or policy performance, while the computational cost grows roughly linearly in $N$. Hence we treat $N$ as a fixed architectural hyperparameter.
Figure 11: Empirical variance of the IQN-based $\mathrm{CVaR}_{0.1}$ estimator as a function of the tail sample size $K$ for the offline-selected RADAC policies on HalfCheetah-Medium-Replay-v2 (left) and Walker2d-Medium-Replay-v2 (right). For each $K\in\{2,4,8,16\}$ we compute $\widehat{\mathrm{CVaR}}_{\alpha}$; the curves show the mean per-state estimator variance with error bars indicating one standard deviation across states for $\alpha\in\{0.05,0.1,0.2\}$; all three risk levels exhibit a similar $1/K$-like decay.
E.3.3 OOD detector robustness
To check that the RADAC << ORAAC trend in Table 2 is not an artifact of the 1-NN score, we repeated the analysis with two additional detectors: a Local Outlier Factor (LOF) and a simple Mahalanobis detector based on a single Gaussian fit to the dataset actions. Table 5 reports the three Stochastic-D4RL medium-expert tasks with these detectors. On HalfCheetah and Walker2d, all three detectors agree with the main-text result and rank RADAC as having substantially lower OOD action rates than ORAAC. On Hopper, RADAC still achieves $5\text{--}8\times$ lower OOD rates under 1-NN and LOF, while we see one exception: the single-Gaussian Mahalanobis detector returns $0\%$ OOD for ORAAC. This reflects a limitation of the detector rather than a reversal of the trend. ORAAC on this task collapses most of its mass onto a single expert-like mode with very small variance, so that all of its actions lie deep inside the $99\%$ confidence ellipsoid of a Gaussian fit to the dataset. A global Mahalanobis score cannot see the thin or multi-modal structure of the dataset manifold and therefore misses the local leakage that is still picked up by the 1-NN and LOF detectors, for which ORAAC continues to have higher OOD rates than RADAC. Aggregating over three seeds, RADAC retains lower 1-NN and LOF OOD rates than ORAAC on all three tasks, and lower Mahalanobis OOD rates on two out of three tasks, indicating that the RADAC << ORAAC ordering is robust to detector. Implementation details of these detectors are detailed in App. F.4
Table 5: Detector–robust OOD action rates $\varepsilon_{\text{act}}$ (%, mean $\pm$ s.e. over 3 seeds) on Stochastic–D4RL medium–expert tasks. Smaller is better.
Env Algo 1-NN LOF Mahalanobis
HalfCheetah-m.e. RADAC 2.01 $\pm$ 0.31 4.42 $\pm$ 0.94 0.09 $\pm$ 0.02
HalfCheetah-m.e. ORAAC 65.24 $\pm$ 4.54 10.02 $\pm$ 1.72 0.34 $\pm$ 0.33
Hopper-m.e. RADAC 0.68 $\pm$ 0.06 2.09 $\pm$ 0.21 0.43 $\pm$ 0.03
Hopper-m.e. ORAAC 3.71 $\pm$ 2.52 14.23 $\pm$ 1.45 0.00 $\pm$ 0.00
Walker2d-m.e. RADAC 1.47 $\pm$ 0.42 0.74 $\pm$ 0.13 0.50 $\pm$ 0.12
Walker2d-m.e. ORAAC 2.80 $\pm$ 0.50 5.81 $\pm$ 1.46 0.76 $\pm$ 0.26
E.4 Runtime and Inference Latency
We compare wall-clock training time and per-action inference latency for DiffusionQL, FlowQL, and RADAC on Hopper-Medium-Expert-v2. All methods use the same A6000 GPU, PyTorch/CUDA stack, and training budget. Training time is measured between the first and last gradient update; latency is the average GPU time to produce a single action for $10^{4}$ replay states with torch.no_grad.
Figure 12: Wall-clock training time (left) and per-action inference latency (right) on Hopper-Medium-Expert-v2.
RADAC replaces the scalar critic with a distributional IQN and optimizes a CVaR objective, but this does not dominate runtime. In these methods, most of the computational cost comes from the expressive generative actor (diffusion or flow), not from the critic. The extra work required by IQN, evaluating a small number of quantiles and averaging the bottom $\alpha$-fraction for CVaR, adds only minor overhead relative to a full diffusion / flow pass.
Conversely, FlowQL carries its own overhead by training both a flow-matching prior and a distilled one-step policy head, and DiffusionQL implementation uses slightly heavier hyperparameters (e.g. a larger actor) than RADAC. As a result, under our implementation and hyperparameter choices, RADAC ends up slightly faster in wall-clock time on Hopper-Medium-Expert-v2. We do not claim that RADAC is intrinsically faster than DiffusionQL or FlowQL; these measurements simply show that the CVaR + distributional critic extension does not introduce an order-of-magnitude runtime penalty. Inference latency is likewise dominated by the shared diffusion/flow backbone, and the RADAC actor achieves comparable or lower per-action latency than the risk-neutral expressive baselines when using similar numbers of denoising / flow steps.
E.5 Return distributions of rollout trajectories
We here visualize the empirical return distributions under the stochastic hazard wrapper for a HalfCheetah-Medium-Expert-v2. For each method we aggregate 30306060 evaluation episodes across three seeds and plot histograms and kernel-density estimates of the rollout returns, together with vertical markers for the mean, median, and $\mathrm{CVaR}_{0.1}$ (Fig. 13). On HalfCheetah-Medium-Expert-v2, for example, RADAC concentrates mass in a high-return band ($\mathrm{CVaR}_{0.1}\!\approx\!4.7\!\times\!10^{3}$) while ORAAC exhibits a multi-modal distribution with occasional near-zero or negative episodes, and DiffusionQL/CQL collapse near zero. Since hazard events in our wrapper correspond to large negative returns and early termination, catastrophic hazard-inducing episodes populate the extreme left tail; the reduced tail mass for RADAC thus reflects a lower frequency of hazard-heavy trajectories rather than mere over-conservatism.
Figure 13: Empirical return distributions on HalfCheetah-Medium-Expert-v2 under the stochastic hazard wrapper. Each subplot shows a histogram and KDE of episodic returns across evaluation rollouts, with vertical lines for the mean, median, and $\mathrm{CVaR}_{0.1}$.
E.6 Risk-return frontier under CVaR level
To make clearer view of the safety/return trade-off, we visualize how the CVaR level $\alpha$ affects RADAC’s behavior on HalfCheetah-Medium-Replay-v2 and Walker2d-Medium-Replay-v2. For each environment, we train RADAC with $\alpha\in\{0.05,0.10,0.20\}$ using the same hyperparameters as in the main Stochastic-D4RL experiments, and aggregate results across multiple seeds. For each run, we select the checkpoint with the highest normalized score and compute the mean normalized return and empirical $\mathrm{CVaR}_{\alpha}$ from eval rollouts under the risky wrapper. We then report the seed-averaged mean normalized score versus $\mathrm{CVaR}_{\alpha}$, with error bars denoting the standard error of the mean across seeds (Fig. 14).
On HalfCheetah-Medium-Replay-v2, varying $\alpha$ across $\{0.05,0.10,0.20\}$ primarily affects the tail: the points move noticeably along the horizontal axis in $\mathrm{CVaR}_{\alpha}$, while the mean normalized score stays in a very tight band, indicating that RADAC can reshape the lower tail of the return distribution with only a minor impact on average performance. In contrast, on Walker2d-Medium-Replay-v2 the frontier moves up and to the right as we adjust $\alpha$: the more tail-sensitive settings simultaneously improve $\mathrm{CVaR}_{\alpha}$ and the mean normalized score, suggesting that on this task catastrophic low-return trajectories are frequent enough that suppressing them not only reduces tail risk but also raises average returns. Overall, these frontiers confirm that the BC+CVaR objective in RADAC exposes a smooth knob to trade off tail-risk and mean return, and that in some regimes increasing tail sensitivity can strictly improve both safety and performance.
Figure 14: Risk-return frontier for RADAC under different CVaR levels. Each point shows the seed-averaged mean normalized score and $\mathrm{CVaR}_{\alpha}$ for RADAC trained with a fixed $\alpha\in\{0.05,0.10,0.20\}$ on HalfCheetah-Medium-Replay-v2 (left) and Walker2d-Medium-Replay-v2 (right). Error bars denote the standard error of the mean across seeds.
Appendix F Experimental Details
F.1 2D Synthetic Task Details
@@ Appendix F Experimental Details
F.2 Stochastic-D4RL MuJoCo Suite
Datasets We adopt the stochastic MuJoCo protocol for risk-sensitive offline RL, following [urpi2021risk]. Policies are evaluated on
$\{\textsc{Hopper},\ \textsc{Walker2d},\ \textsc{HalfcCheetah}\}\ \times\ \{\textsc{medium-expert},\ \textsc{medium-replay}\},$
Compared to prior work, we prefer medium-expert and medium-replay to validate both risk sensitivity and policy expressiveness under multimodal action distributions. For training, we relabel per-transition rewards in the offline datasets to inject stochastic hazards (velocity or torso-pitch thresholds with Bernoulli penalties and early termination); the same hazard model is used at evaluation. All algorithms (CQL, CODAC, ORAAC, DiffusionQL/FlowQL, RADAC, etc.) are trained on these relabeled rewards; the hazard indicator is never provided as an input feature or mask, so no method receives privileged information about hazard locations. This ensures the critic and the policy are trained on the risk-aware rewards rather than only being tested under hazards.
@@ F.2 Stochastic-D4RL MuJoCo Suite
F.3 Baselines: Implementation & Hyperparameters
@@
  • CQL [kumar2020conservative] (value pessimism). Non-distributional conservative Q-learning baseline.
  • CODAC [ma2021conservative] (distributional conservative learning). We primarily use the CVaR-optimizing specification (“CODAC-C”, $\mathrm{CVaR}_{0.1}$ objective).
  • ORAAC [urpi2021risk] (offline risk-averse actor–critic). A distributional critic with imitation-regularized policy optimizing a coherent risk objective.
  • DiffusionQL [wang2022diffusion] (expressive risk-neutral diffusion policy).
  • FlowQL [park2025flow] (expressive risk-neutral flow-matching policy).
@@
For each of baselines, we run all baselines ourselves and tune the following parameters or adopt authors’ recommended settings, mirroring the practice in [kumar2020conservative; ma2021conservative; park2025flow; urpi2021risk; wang2022diffusion].
@@
  • FlowQL [park2025flow]: we sweep the policy weight $\alpha\in\{1,10,30,100,1000\}$ per task and report the best-performing setting (selection by $\mathrm{CVaR}_{0.1}$ unless noted).
@@
  • DiffusionQL [wang2022diffusion]: we consider $\eta\in\{0.1,0.5,1.0\}$ for BC coefficient . we use authors’ recommended configuration for other parameters without retuning. We also used the best checkpoint of their model on each benchmark by following their protocol.
@@
  • ORAAC [urpi2021risk]: use the paper’s recommended configuration (distributional critic, risk level $\alpha=0.1$, anchor/prior regularization) without additional sweeps.
@@
  • CODAC [ma2021conservative]: use the paper’s tuned settings for D4RL (risk level $\alpha=0.1$) without further tuning.
@@
  • CQL [kumar2020conservative]: use the standard conservative coefficient and implementation defaults for MuJoCo locomotion.
F.4 Estimating OOD action rates and detectors
@@
At evaluation time we measure the fraction of actions produced by a policy that fall outside the empirical action support of the offline dataset. Let $\mathcal{A}_{\mathcal{D}}=\{a_{i}\}_{i=1}^{N}$ denote the set of offline actions for a given task, and let $\mathcal{A}_{\text{eval}}=\{a^{(\text{eval})}_{t}\}_{t=1}^{T}$ be all actions emitted across evaluation rollouts (we use $S$ seeds and $10$ episodes per seed; S=5S{=}5 in the main table and S=3S{=}3 in the detector-robustness ablation). Actions are already scaled to [1,1][-1,1] per dimension in MuJoCo, so we work directly in $\ell_{2}$ action space.
Given a detector that assigns an OOD indicator
$\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})\in\{0,1\},$
we define the OOD action rate
$\varepsilon_{\text{act}}\;=\;\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})$,
@@
and report the mean and standard error over seeds. Because episodes may terminate early under the stochastic wrappers, $T$ is the actual number of executed timesteps, which makes rates comparable across seeds.
1-NN detector (main text).
For each dataset action $a_{i}$ we compute its nearest neighbour among the other dataset actions,
$d_{i}\;=\;\min_{j\neq i}\,\|a_{i}-a_{j}\|_{2},\quad\operatorname{medNN}\;=\;\mathrm{median}\{d_{i}\}_{i=1}^{N},$
and set the OOD threshold $\tau=\kappa\cdot\operatorname{medNN}$ with $\kappa{=}3$. For each evaluation action,
$d^{(\text{eval})}_{t}\;=\;\min_{i}\,\|a^{(\text{eval})}_{t}-a_{i}\|_{2},\quad\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t})=\mathbb{I}\!\{d^{(\text{eval})}_{t}>\tau\}.$
Distances are computed efficiently via a KD–tree built on $\mathcal{A}_{\mathcal{D}}$. This 1–NN-based $\varepsilon_{\text{act}}$ is the quantity reported in Sec. 5.3.
Alternative detectors.
To check that the RADAC << ORAAC trend is not an artifact of the 1–NN score, we also evaluate two density- and neighbourhood-based detectors on the same action space:
  • Local Outlier Factor (LOF). We use the standard LOF score with $20$ neighbours and contamination level $0.01$. The detector is trained on $\mathcal{A}_{\mathcal{D}}$, and $\mathbf{1}_{\text{OOD}}(a^{(\text{eval})}_{t}){=}1$ when LOF predicts an outlier label for $a^{(\text{eval})}_{t}$.
  • Single-Gaussian Mahalanobis distance. We fit a single multivariate Gaussian $\mathcal{N}(\mu,\Sigma)\,$ to $\mathcal{A}_{\mathcal{D}}$, with a small diagonal jitter added to $\Sigma$ for numerical stability. For any action $a$ we compute the squared Mahalanobis distance
    $d^{2}_{\mathrm{Mah}}(a)\;=\;(a-\mu)^{\top}\Sigma^{-1}(a-\mu).$
    We estimate an upper-threshold $\tau$ as the $99\%$ quantile of $\{d^{2}_{\mathrm{Mah}}(a_{i})\}_{i}$ using up to $5\!\times\!10^{4}$ subsampled dataset actions, and declare $a^{(\text{eval})}_{t}$ OOD if $d^{2}_{\mathrm{Mah}}(a^{(\text{eval})}_{t})>\tau$.
@@
For a seed-level rate $\hat{\varepsilon}$ computed by any of the detectors above, we approximate the standard error via a binomial model, $\mathrm{SE}=\sqrt{\hat{\varepsilon}(1-\hat{\varepsilon})/T},$ and report the across-seed mean $\pm$ s.e.